Typist A spends 2560 minutes to do a job. How much time does it take for Typist B to do the same job, if when they work together it takes them both 45 minutes to do four jobs?
I am having trouble with this thanks

Answer by mananth(16946)

I am not too sure the problem has been copied correctly.
Anyway
45 minutes to do 4 jobs
1 job 45/4 = minutes.
together they do 4/45 of the job in 1 minute.
...
Typist A takes 2560 minutes to do the job
he does 1/2560 of the job in 1 minute.
...
Typist B does 4/45 - 1/2650 of the job in 1 minute
4*2650-45 / 45*2650
=10600-45/119250
=10555/119250
=0.089 of the job in 1 minute
Typist B will take 1/0.089 minutes
11.3 minutes